# CMAT: A Multi-Agent Collaboration Tuning Framework for Enhancing Small Language Models

### v4

@@ Abstract
Open large language models (LLMs) have significantly advanced the field of natural language processing, showcasing impressive performance across various tasks. Despite the significant advancements in LLMs, their effective operation still relies heavily on human input to accurately guide the dialogue flow, with agent tuning being a crucial optimization technique that involves human adjustments to the model for better response to such guidance. Addressing this dependency, our work introduces the TinyAgent model, trained on a meticulously curated high-quality dataset. We also present the Collaborative Multi-Agent Tuning (CMAT) framework, an innovative system designed to augment language agent capabilities through adaptive weight updates based on environmental feedback. This framework fosters collaborative learning and real-time adaptation among multiple intelligent agents, enhancing their context-awareness and long-term memory. In this research, we propose a new communication agent framework that integrates multi-agent systems with environmental feedback mechanisms, offering a scalable method to explore cooperative behaviors. Notably, our TinyAgent-7B model exhibits performance on par with GPT-3.5, despite having fewer parameters, signifying a substantial improvement in the efficiency and effectiveness of LLMs 111Access models, demos at https://github.com/heimy2000/CMAT/tree/master.
1 Introduction
In the rapid development of the field of artificial intelligence, large language models (LLMs) such as BERT and GPT-4 OpenAI (2023) have become important cornerstones of natural language processing (NLP). These models utilize the Transformer architecture and effectively capture long-distance dependencies through multi-head self-attention mechanisms, demonstrating strong capabilities across various NLP tasks. With technological advancements, the performance and application scope of LLMs continue to expand, promising significant improvements in computational efficiency and functionality, including anticipated advanced features such as self-improvement, self-checking, and sparse expert models Liu et al. (2023).
@@ 1 Introduction
Our research is driven by the need to overcome the significant challenges faced by LLMs in real-world deployments, particularly the high computational resource requirements, data biases, and lack of robustness. These issues limit their applicability in resource-constrained environments and highlight the urgency of enhancing model efficiency and adaptability Abid et al. (2021); Du et al. (2022). As demonstrated by Figure 1, we aim to address these limitations by optimizing models and training methods to enable smaller models to match the performance levels of larger models. Additionally, recognizing the potential of MAS to improve processing efficiency and system adaptability through agent cooperation, we seek to develop a collaborative agent framework. This framework aims to facilitate effective cooperation among agents, thereby overcoming the performance gap and propelling further research and innovation in the field of LLMs Ferry et al. (2018); Talwar et al. (2005). In our experiments, we evaluated the capabilities of large models with and without the use of prompts and observed that low-quality prompts can significantly degrade model performance. Consequently, we propose the Collaborative Multi-Agent Tuning (CMAT) framework.
The CMAT framework introduces a structured environment where individual agents, each with specialized roles and capabilities, work together to process information, make decisions, and solve complex tasks Hernández-Orallo et al. (2017). By sharing insights and learning from interactions within this multi-agent ecosystem, the framework allows for a more scalable and flexible approach to training LLMs Lewis et al. (2017). This collaborative effort not only helps in bridging the gap in performance between smaller and larger models but also fosters a more resilient system capable of adapting to new challenges without extensive human intervention Kaplan et al. (2020). Through CMAT, we aim to push the boundaries of what is possible with LLMs, making them more accessible and effective for a wider range of applications Rajpurkar et al. (2018).
@@
We evaluated the fine-tuned TinyAgent models across multiple agent tasks, finding that in certain scenarios, their performance rivals that of advanced LLMs like GPT-4 and agentlm Zeng et al. (2023), demonstrating the potential efficiency and capabilities of compact models.
2 Related Work
2.1 LLMs Applications in a Multi-Agent Framework
@@
We explore the applications of LLMs within multi-agent systems, highlighting their role versatility as users, assistants, and checkers, and their capability to offer bespoke support and solutions across such environments de Zarzà et al. (2023); Talebirad and Nadiri (2023). LLMs showcase remarkable adaptability to tasks through methods like supervised fine-tuning and real-time feedback learning, notably in tasks that require a sophisticated understanding and execution related to operating systems or databases Christianos et al. (2023); Li et al. (2023). Furthermore, LLMs are adept at enhancing communication and collaboration among agents, a critical component for addressing complex issues that necessitate multi-role coordination Zhao et al. (2021). Nevertheless, LLMs encounter specific challenges within multi-agent frameworks, especially in situations that demand a nuanced contextual comprehension and sustained memory retention, as well as adapting to fast-evolving environments and unforeseeable tasks Diallo et al. (2020). Issues such as data bias, security concerns, and the intricacies of crafting effective protocols for multi-agent cooperation stand as significant hurdles in this domain Zhang et al. (2017); García et al. (2015). Thus, by summarizing LLMs’ roles in multi-agent frameworks, we underscore the critical need for continued innovation and research exploration, aimed at overcoming these technological hurdles and leveraging the full potential of LLMs in complex systems Lu and Zhang (2020).
@@
To enhance the adaptability and collaborative capabilities of LLMs in multi-agent systems, we’ve implemented memory modes, including long-term support and short-term memory with environmental feedback Liang et al. (2016). This allows LLMs to better interact, learn, and adapt in dynamic environments, leveraging past experiences and responding to changes swiftly.
2.2 The tuning method for LLMs
@@
The main tuning methods include supervised fine-tuning and reinforcement learning Ouyang et al. (2022). Supervised fine-tuning enhances performance by training models on specific task datasets, and is especially suitable for tasks such as natural language understanding (NLU) Howard and Ruder (2018). On the other hand, reinforcement learning, guided by reward mechanisms, is suitable for handling complex and variable tasks Mnih et al. (2015). The effective combination of these two methods can significantly improve the performance of LLMs in various tasks. Notably, LLMs of reduced scale, such as those encompassing 1.8 billion parameters, can achieve performance levels akin to those of models with greater parameter counts, like 6 billion parameters, when supported by high-quality datasets Stiennon et al. (2020). This demonstrates that excellent data quality and appropriate tuning strategies play a decisive role in the performance of LLMs. Therefore, investing efforts in improving data quality and choosing the right tuning methods is essential for achieving optimal performance of LLMs in various application scenarios Howard and Ruder (2018). Through our work combining supervised fine-tuning with reinforcement learning, we’ve notably advanced LLM performance across a spectrum of tasks, showcasing significant improvements in task-specific benchmarks Ouyang et al. (2022).
3 Methodology
Our research is dedicated to the development and application of the Collaborative Multi-Agent Language Model Tuning (CMAT) framework. This framework aims to enhance the quality of decision-making, controllability, and operational efficiency within complex systems through the collaborative efforts of various agent roles, including the User, Assistant, and Checker. The CMAT framework employs advanced supervised fine-tuning techniques such as Low-Rank Adaptation (LoRA)Hu et al. (2021) and Prompt Tuning (PTuning)Lester et al. (2021), leveraging pre-trained datasets like AgentBench. Additionally, the framework incorporates interactive environment engagement and memory updates to further refine the model’s performance, drawing inspiration from the principles of Reinforcement Learning from Human Feedback (RLHF)Vázquez-Canteli and Nagy (2019).
3.1 Actor-Evaluator Dynamics in the CMAT Framework
In multi-agent systems, the Actor plays a crucial role in behavior generation through LLMs. The Actor responds to environmental state observations to generate text and actions, while also using reflection to assess and adjust its behavior Wang and Shi (2019). In traditional reinforcement learning (RL) frameworks, the Actor decides on actions based on a specific policy and receives information about the current state of the environment. In the CMAT framework, the Actor model utilizes approaches like Chain of Thought and ReAct, enhancing the traditional text generation models to delve into different facets of behavior generation.
3.2 Real-Time Review Loop and Interdisciplinary Research
Long-term Memory
Within the context of multi-agent systems, the significance of long-term memory for LLMs cannot be overstatedKato et al. (2022). It serves not merely as a storage for information but as an integrated knowledge management system, enabling LLMs to retain and access vital data over extended periodsBhadra (2022). Our research delves into interdisciplinary insights, particularly from psycholinguistics, to enhance the problem-solving capabilities of LLMs through the strategic utilization of Chain-of-Thought (CoT).
A pivotal discovery from our interdisciplinary exploration is the impact of CoT generation order on the correctness of LLMs’ responses. We have observed that when LLMs first generate a CoT before producing an answer, the correctness of the final answer is significantly higher compared to when the answer is generated first and then a CoT is constructed. Notably, the CoTs generated in both scenarios are identical, indicating that the order of generation is crucial. This suggests that while the models may have the potential to arrive at correct conclusions, their capabilities are not yet sufficient to do so autonomously. By employing a memory-assisted approach that prioritizes CoT generation, we can effectively mitigate this limitation and enhance the overall performance of LLMs in complex decision-making tasks.
Short-term Memory and Environmental Feedback
Conversely, short-term memory focuses on immediate contexts and the rapid processing of new information, which is essential for LLMs to adapt swiftly to dynamic environmentsPae et al. (2012); Liu and Guo (2019). This facet of memory is equally informed by psycholinguistic research, enabling LLMs to respond promptly to environmental changes and demands. Furthermore, environmental communication feedback mechanisms complement short-term memory by providing real-time inputs and interactions within the multi-agent frameworkYogatama et al. (2021). This continuous flow of information allows LLMs to adjust their strategies based on the latest data, facilitating a more agile and effective adaptation process.
By integrating the strengths of both long-term and short-term memory, informed by interdisciplinary research, our LLMs are better positioned to simulate human cognitive processes, offering a more human-like interaction experience. This cross-disciplinary approach not only deepens our understanding of memory system design but also paves the way for future AI research, optimizing LLMs’ performance and adaptability through the emulation of human thought processes.
@@ 3.2 Real-Time Review Loop and Interdisciplinary Research
The Reflexion Process
Artificial intelligence systems are increasingly adopting self-reflection mechanisms to enhance their decision-making capabilities. By systematically reviewing past actions and their outcomes, these systems can identify patterns and make informed adjustments to their strategies.
The CMAT framework introduces a Reflexion Process of collaboration between computation and cognition by initializing LLMs, users, assistants, checkers, and a set of tasks Li et al. (2018). Each task is executed by an assistant and verified by a checker. The assistant selects an action through interaction with LLMs, obtaining a new state and reward Che et al. (2021). If the checker verifies as correct, the system updates the state to and stores in memory mem; otherwise, it adjusts strategy and retries. The entire process not only involves confirmation of correct actions but also includes updating the strategy based on feedback from the checker, creating a continuous cycle of learning and adjustment. Each iteration aims to optimize the assistant’s decision-making strategy through practice and reflection, enhancing the efficiency and accuracy of task execution. Once all tasks are verified, it outputs the final states and evaluation results of all tasks Silver et al. (2017).Within this cycle, there is also a critical step: updating the checker’s strategy. This means that after each update of the strategy , not only does the assistant’s strategy need to be adjusted based on feedback, but the checker’s verification strategy also needs to be updated synchronously to ensure its decision-making logic is consistent with the latest environmental state and strategy. Such updates ensure the checker can accurately assess the assistant’s actions, further enhancing the system’s overall performance and reliability. Through this approach, the CMAT framework achieves a deeper level of collaboration and self-optimization between computation and cognition, enabling AI systems to face new challenges and environments with higher efficiency and adaptability.
Checker-In-The-Loop
To enhance the controllability of the role-playing framework, we have introduced an inspector agent that can inspect the role-playing Tao et al. (2023) agents for any issues. This enables a decision-making process for task solving similar to tree search. In practice, the inspector can be either an AI agent or a human.
3.3 Compared to Existing Practice
Our method stands out by dynamically updating memory through real-time interactions, enabling rapid adaptation to new experiences. It uses direct feedback from the environment for quick learning and incorporates both short-term and long-term memory updates for efficient decision-making. Our approach, enhanced with self-reflection and experience replay, offers deeper understanding and better response to complex scenarios, leading to more precise and rational decisions in changing conditions.
@@ 3.3 Compared to Existing Practice
| chatglm2-6b 222https://github.com/thudm/chatglm2-6b | v1.1 | 4.2 | 1.3 | 0.0 | 0.0 | 0.0 | 0.0 | |
@@
| llama2-7b Touvron et al. (2023) | chat | 0.0 | 4.2 | 8.0 | 0.0 | 11.6 | 7.0 | |
| zephyr-7b Tunstall et al. (2023) | alpha | 12.5 | 9.7 | 5.0 | 8.0 | 45.0 | 11.0 | |
| baichuan2-6b Yang et al. (2023) | chat | 2.8 | 9.7 | 0.0 | 0.0 | 6.1 | 11.0 | |
| mpt-7b 333https://github.com/mosaicml/llm-foundry/ | chat | 5.6 | 9.7 | 12.7 | 0.0 | 0.0 | 0.0 | |
@@
4 Experiments
Our evaluation framework rigorously tests intelligent agents in six key domains to ensure their readiness for diverse real-world challenges Ross et al. (2023). These areas include seamless LLM integration into OS with an emphasis on security and user interaction; proficiency in real DB operations using SQL Halevy et al. (2004); task execution on the simulated e-commerce platform WebShop(WS) Yao et al. (2022); constructing and using KGs for enhanced semantic understanding; employing the M2W dataset for complex web tasks, marking the first dataset for developing general web agents following language instructions; and applying abstract reasoning and visual tasks in the text-based ALFWorld(ALF) Shridhar et al. (2021). For more implementation and evaluation details, see Appendices A and B.
4.1 Dataset
The dataset for our research was meticulously constructed to comprehensively evaluate the capabilities of agents Gou et al. (2020). It was established through self-collected methods, aimed at providing a rich and diverse testing environment to thoroughly assess the performance of deep learning models across various tasks Sachdeva and McAuley (2023). The construction of the dataset included key processes such as data collection, filtering, enhancement, and knowledge distillation Chen and Liu (2018). Through detailed screening and processing, we ensured the accuracy and consistency of the dataset, retaining only high-quality samples directly related to the testing objectives Sachdeva and McAuley (2023). Faced with issues of data imbalance and insufficient samples, we utilized data augmentation and knowledge distillation techniques. Knowledge distillation helped us to extract the most valuable and representative information from the vast amount of collected data, thus building an efficient and refined testing dataset. This process significantly improved the quality and applicability of the dataset, providing a solid foundation for evaluating the capabilities of model agents Mishra and Marr (2017).
4.2 Evaluating Code Correction
As shown in the Table 1, in this study, we conducted a comprehensive performance evaluation of TinyAgent-1.8B and the CodeLlama series models (CodeLlama7B and CodeLlama13B), aiming to explore their multi-task checking capabilities, including but not limited to code correction, OS configuration, DB query optimization, and WS. The experimental results showed that TinyAgent-1.8B demonstrated a significant advantage in cross-task performance evaluation compared to the CodeLlama series models. This performance was not only significant in code correction tasks but also prominent in other checking tasks such as OS configuration, DB query optimization, and WS management. These findings highlight that TinyAgent-1.8B not only possesses efficient code analysis capabilities but is also widely applicable to the inspection and optimization of other complex systems.
4.3 Baselines
| Models | OS | DB | KG | ALF | WS | M2W |
| tinyagent-7b | 27.3 | 43.0 | 38.0 | 10.0 | 61.8 | 14.0 |
| - agent only | 20.1 | 39.3 | 25.0 | 2.0 | 55.7 | 7.0 |
| - general only | 9.7 | 5.4 | 0.0 | 0.0 | 26.6 | 5.0 |
4.4 Results analysis

### v6

@@ Abstract
Open large language models (LLMs) have significantly advanced the field of natural language processing, showcasing impressive performance across various tasks. Despite the significant advancements in LLMs, their effective operation still relies heavily on human input to accurately guide the dialogue flow, with agent tuning being a crucial optimization technique that involves human adjustments to the model for better response to such guidance. Addressing this dependency, our work introduces the TinyAgent model, trained on a meticulously curated high-quality dataset. We also present the Collaborative Multi-Agent Tuning (CMAT) framework, an innovative system designed to augment language agent capabilities through adaptive weight updates based on environmental feedback. This framework fosters collaborative learning and real-time adaptation among multiple intelligent agents, enhancing their context-awareness and long-term memory. In this research, we propose a new communication agent framework that integrates multi-agent systems with environmental feedback mechanisms, offering a scalable method to explore cooperative behaviors. Notably, our TinyAgent-7B model exhibits performance on par with GPT-3.5, despite having fewer parameters, signifying a substantial improvement in the efficiency and effectiveness of LLMs.
Index Terms:
Collaborative Multi-Agent Framework, Language Model Fine-Tuning, Reinforcement Learning from Human FeedbackI Introduction
In the rapid development of the field of artificial intelligence, large language models (LLMs) such as BERT and GPT-4 [1] have become important cornerstones of natural language processing (NLP). These models utilize the Transformer architecture and effectively capture long-distance dependencies through multi-head self-attention mechanisms, demonstrating strong capabilities across various NLP tasks. With technological advancements, the performance and application scope of LLMs continue to expand, promising significant improvements in computational efficiency and functionality, including anticipated advanced features such as self-improvement, self-checking, and sparse expert models [2].
@@
Our research is driven by the need to overcome the significant challenges faced by LLMs in real-world deployments, particularly the high computational resource requirements, data biases, and lack of robustness. These issues limit their applicability in resource-constrained environments and highlight the urgency of enhancing model efficiency and adaptability [5, 6]. As demonstrated by Figure 1, we aim to address these limitations by optimizing models and training methods to enable smaller models to match the performance levels of larger models. Additionally, recognizing the potential of MAS to improve processing efficiency and system adaptability through agent cooperation, we seek to develop a collaborative agent framework. This framework aims to facilitate effective cooperation among agents, thereby overcoming the performance gap and propelling further research and innovation in the field of LLMs [7, 8]. In our experiments, we evaluated the capabilities of large models with and without the use of prompts and observed that low-quality prompts can significantly degrade model performance. Consequently, we propose the Collaborative Multi-Agent Tuning (CMAT) framework.
The CMAT framework introduces a structured environment where individual agents, each with specialized roles and capabilities, work together to process information, make decisions, and solve complex tasks [9]. By sharing insights and learning from interactions within this multi-agent ecosystem, the framework allows for a more scalable and flexible approach to training LLMs [10]. This collaborative effort not only helps in bridging the gap in performance between smaller and larger models but also fosters a more resilient system capable of adapting to new challenges without extensive human intervention [11]. Through CMAT, we aim to push the boundaries of what is possible with LLMs, making them more accessible and effective for a wider range of applications [12].
@@
We evaluated the fine-tuned TinyAgent models across multiple agent tasks, finding that in certain scenarios, their performance rivals that of advanced LLMs like GPT-4 and agentlm [13], demonstrating the potential efficiency and capabilities of compact models.
II Related Work
II-A LLMs Applications in a Multi-Agent Framework
@@
We explore the applications of LLMs within multi-agent systems, highlighting their role versatility as users, assistants, and checkers, and their capability to offer bespoke support and solutions across such environments [14, 15]. LLMs showcase remarkable adaptability to tasks through methods like supervised fine-tuning and real-time feedback learning, notably in tasks that require a sophisticated understanding and execution related to operating systems or databases [16, 17]. Furthermore, LLMs are adept at enhancing communication and collaboration among agents, a critical component for addressing complex issues that necessitate multi-role coordination [18]. Nevertheless, LLMs encounter specific challenges within multi-agent frameworks, especially in situations that demand a nuanced contextual comprehension and sustained memory retention, as well as adapting to fast-evolving environments and unforeseeable tasks [19]. Issues such as data bias, security concerns, and the intricacies of crafting effective protocols for multi-agent cooperation stand as significant hurdles in this domain [20, 21]. Thus, by summarizing LLMs’ roles in multi-agent frameworks, we underscore the critical need for continued innovation and research exploration, aimed at overcoming these technological hurdles and leveraging the full potential of LLMs in complex systems [22].
@@
To enhance the adaptability and collaborative capabilities of LLMs in multi-agent systems, we’ve implemented memory modes, including long-term support and short-term memory with environmental feedback [23]. This allows LLMs to better interact, learn, and adapt in dynamic environments, leveraging past experiences and responding to changes swiftly.
II-B The tuning method for LLMs
@@
The main tuning methods include supervised fine-tuning and reinforcement learning [24]. Supervised fine-tuning enhances performance by training models on specific task datasets, and is especially suitable for tasks such as natural language understanding (NLU) [25]. On the other hand, reinforcement learning, guided by reward mechanisms, is suitable for handling complex and variable tasks [26]. The effective combination of these two methods can significantly improve the performance of LLMs in various tasks. Notably, LLMs of reduced scale, such as those encompassing 1.8 billion parameters, can achieve performance levels akin to those of models with greater parameter counts, like 6 billion parameters, when supported by high-quality datasets [27]. This demonstrates that excellent data quality and appropriate tuning strategies play a decisive role in the performance of LLMs. Therefore, investing efforts in improving data quality and choosing the right tuning methods is essential for achieving optimal performance of LLMs in various application scenarios [25]. Through our work combining supervised fine-tuning with reinforcement learning, we’ve notably advanced LLM performance across a spectrum of tasks, showcasing significant improvements in task-specific benchmarks [24].
III Methodology
The Collaborative Multi-Agent Language Model Tuning (CMAT) framework improves decision-making, controllability, and efficiency in complex systems by coordinating three roles: User (), Assistant (), and Checker (). We fine-tune language models using LoRA [28], P-Tuning [29], and integrate ideas from RLHF [30], Chain of Thought (CoT), and ReAct to enhance reasoning and adaptability.
III-A Agent Roles and Actor-Critic Dynamics
In CMAT, the User () provides inputs, the Assistant () acts as the Actor generating actions, and the Checker () serves as the Critic providing feedback. At time , given input , the Assistant produces action based on its policy . The Checker evaluates this action and returns feedback , guiding continuous policy refinement.
III-B Learning Strategy
III-B1 Supervised Fine-Tuning
We start by fine-tuning the Assistant model on a labeled dataset :
where is the cross-entropy loss. This step ensures the model can produce useful actions before incorporating feedback loops.
III-B2 Incorporating CoT and ReAct
To enhance reasoning, we adopt Chain of Thought (CoT) to generate intermediate reasoning steps before the final action:
Additionally, ReAct interleaves reasoning and acting tokens, encouraging more accurate and context-aware decisions.
Empirically, generating reasoning steps (CoT) first, then the final answer, improves correctness. This suggests that explicitly formulating the thought process guides the model towards better actions.
III-B3 Feedback-Driven Policy Optimization
After supervised fine-tuning, the Assistant refines its policy using Checker feedback. We adopt an Actor-Critic-like scheme without full RL exploration. The Assistant updates its parameters via:
where is the error term:
Here, is derived from feedback . The Checker updates its value function:
Unlike traditional RL that involves environment exploration, here the process relies on direct Checker feedback. The iterative cycle refines the Assistant’s decision-making, ensuring alignment with the desired standards.
III-C Checker-In-The-Loop Mechanism
By positioning the Checker directly in the training loop, we ensure that the Assistant receives immediate corrective signals whenever it deviates from expected behavior
III-D Memory Management and Self-Reflection
The Assistant maintains two memories: short-term and long-term . Short-term memory tracks recent context:
while long-term memory stores significant insights via self-reflection:
Using the reflection, the Assistant updates its parameters:
where encodes feedback-driven loss and incorporates improvements identified via reflection.
IV Experiments
Our evaluation framework rigorously tests intelligent agents in six key domains to ensure their readiness for diverse real-world challenges [31]. These areas include seamless LLM integration into OS with an emphasis on security and user interaction; proficiency in real DB operations using SQL [32]; task execution on the simulated e-commerce platform WebShop(WS) [33]; constructing and using KGs for enhanced semantic understanding; employing the M2W dataset for complex web tasks, marking the first dataset for developing general web agents following language instructions; and applying abstract reasoning and visual tasks in the text-based ALFWorld(ALF) [34].
IV-A Dataset
The dataset for our research was meticulously constructed to comprehensively evaluate the capabilities of agents [35]. It was established through self-collected methods, aimed at providing a rich and diverse testing environment to thoroughly assess the performance of deep learning models across various tasks [36]. The construction of the dataset included key processes such as data collection, filtering, enhancement, and knowledge distillation [37]. Through detailed screening and processing, we ensured the accuracy and consistency of the dataset, retaining only high-quality samples directly related to the testing objectives [36]. Faced with issues of data imbalance and insufficient samples, we utilized data augmentation and knowledge distillation techniques. Knowledge distillation helped us to extract the most valuable and representative information from the vast amount of collected data, thus building an efficient and refined testing dataset. This process significantly improved the quality and applicability of the dataset, providing a solid foundation for evaluating the capabilities of model agents [38].
IV-B Evaluating Code Correction
As shown in the Table I, in this study, we conducted a comprehensive performance evaluation of TinyAgent-1.8B and the CodeLlama series models (CodeLlama7B and CodeLlama13B), aiming to explore their multi-task checking capabilities, including but not limited to code correction, OS configuration, DB query optimization, and WS.
@@ IV-B Evaluating Code Correction
The experimental results showed that TinyAgent-1.8B demonstrated a significant advantage in cross-task performance evaluation compared to the CodeLlama series models. This performance was not only significant in code correction tasks but also prominent in other checking tasks such as OS configuration, DB query optimization, and WS management. These findings highlight that TinyAgent-1.8B not only possesses efficient code analysis capabilities but is also widely applicable to the inspection and optimization of other complex systems.
IV-C Baselines
@@
| chatglm2-6b 111https://github.com/thudm/chatglm2-6b | v1.1 | 4.2 | 1.3 | 0.0 | 0.0 | 0.0 | 0.0 | |
@@
| llama2-7b [41] | chat | 0.0 | 4.2 | 8.0 | 0.0 | 11.6 | 7.0 | |
| zephyr-7b [42] | alpha | 12.5 | 9.7 | 5.0 | 8.0 | 45.0 | 11.0 | |
| baichuan2-6b [43] | chat | 2.8 | 9.7 | 0.0 | 0.0 | 6.1 | 11.0 | |
| mpt-7b 222https://github.com/mosaicml/llm-foundry/ | chat | 5.6 | 9.7 | 12.7 | 0.0 | 0.0 | 0.0 | |
@@
IV-D Results analysis